# Instance-optimal Clipping for Summation Problems in the Shuffle Model of Differential Privacy

### v1

@@ Abstract
Differentially private mechanisms achieving worst-case optimal error bounds (e.g., the classical Laplace mechanism) are well-studied in the literature. However, when typical data are far from the worst case, instance-specific error bounds—which depend on the largest value in the dataset—are more meaningful. For example, consider the sum estimation problem, where each user has an integer $x_{i}$ from the domain $\{0,1,\dots,U\}$ and we wish to estimate $\sum_{i}x_{i}$. This has a worst-case optimal error of $O(U/\varepsilon)$, while recent work has shown that the clipping mechanism can achieve an instance-optimal error of $O(\max_{i}x_{i}\cdot\log\log U/\varepsilon)$. Under the shuffle model, known instance-optimal protocols are less communication-efficient. The clipping mechanism also works in the shuffle model, but requires two rounds: Round one finds the clipping threshold, and round two does the clipping and computes the noisy sum of the clipped data. In this paper, we show how these two seemingly sequential steps can be done simultaneously in one round using just $1+o(1)$ messages per user, while maintaining the instance-optimal error bound. We also extend our technique to the high-dimensional sum estimation problem and sparse vector aggregation (a.k.a. frequency estimation under user-level differential privacy). Our experiments show order-of-magnitude improvements of our protocols in terms of error compared with prior work.
@@ 1 Introduction
The shuffle model [${}^{+}$17, ${}^{+}$19, ${}^{+}$19] of differential privacy (DP) is widely-studied in the context of DP computation over distributed data. The model has 3 steps: (1) Each client uses a randomizer $\mathcal{R}(\cdot)$ to privatize their data. (2) A trusted shuffler randomly permutes the inputs from each client and passes them to an untrusted analyzer, which (3) conducts further analysis. Unlike the central model of DP, where a trusted curator has access to all the data, the shuffle model provides stronger privacy protection by removing the dependency on a trusted curator. Unlike the local model, where clients send noisy results to the analyzer directly, the addition of the trusted shuffler allows for a significantly improved privacy-accuracy tradeoff. For problems like bit counting, shuffle-DP achieves an error of $O(1/\varepsilon)$ with constant probability [^1] In Section 1, all stated error guarantees hold with constant probability. We will make the dependency on the failure probability $\beta$ more explicit in later sections. [GKMP20], matching the best error of central-DP, while the smallest error achievable under local-DP is $O(\sqrt{n}/\varepsilon)$ [${}^{+}$17, DNOR16].
@@
The summation problem, a fundamental problem with applications in statistics [KLSU19, BDKU20, HLY21], data analytics [${}^{+}$22, THMR20, DSY23], and machine learning such as the training of deep learning models [SCS13, BST14, ${}^{+}$16, ${}^{+}$18] and clustering algorithms [Ste21, SK18], has been studied in many works under the shuffle model [BBGN19, ${}^{+}$19, GPV19, GMPV20, BBGN20, ${}^{+}$21]. In this problem, each user $i\in[n]:=\{1,\dots,n\}$ holds an integer $x_{i}\in\{0,1,\dots,U\}$ and the goal is to estimate $\mathrm{Sum}(D):=\sum_{i}x_{i}$, where $D=(x_{1},\dots,x_{n})$. All of these works for sum estimation under shuffle-DP focus on achieving an error of $O(U/\varepsilon)$. Such an error can be easily achieved under central-DP, where the curator releases the true sum after masking it with a Laplace noise of scale $U/\varepsilon$. In the shuffle-DP model, besides error, another criterion that should be considered is the communication cost. The recent shuffle-DP summation protocol of [${}^{+}$21] both matches the error of central-DP and achieves optimal communication. More precisely, it achieves an error that is just $1+o(1)$ times that of the Laplace mechanism, while each user just sends in expectation $1+o(1)$ messages, each of a logarithmic number of bits.
@@
However, in real applications (as well as in [${}^{+}$21]), $U$ must be set independently of the dataset; to account for all possible datasets, it should be conservatively large. For instance, if we only know that the $x_{i}$’s are 32-bit integers, then $U=2^{32}-1$. Then the error of $O(U/\varepsilon)$ could dwarf the true sum for most datasets. Notice that sometimes some prior knowledge is available, and then a smaller $U$ could be used. For example, if we know that the $x_{i}$’s are people’s incomes, then we may set $U$ as that of the richest person in the world. Such a $U$ is still too large for most datasets as such a rich person seldom appears in most datasets.
@@
Instance-Awareness.  The earlier error bound of $O(U/\varepsilon)$ can be shown to be optimal, but only in the worst case. When typical input data are much smaller than $U$, an instance-specific mechanism (and error bound) can be obtained—i.e., a mechanism whose error depends on the largest element of the dataset. In the example of incomes above, an instance-aware mechanism would achieve an error proportional to the actual maximum income in the dataset. This insight has recently been explored under the central model of DP [ATMR21, ${}^{+}$19, MRTZ17, HLY21, FDY22, ${}^{+}$22].
@@
A widely used technique for achieving instance-specific error bounds under central-DP is the clipping mechanism [ATMR21, ${}^{+}$19, MRTZ17, HLY21]. For some $\tau$, each $x_{i}$ is clipped to $\mathrm{Clip}(x_{i},\tau):=\min(x_{i},\tau)$. Then we compute the sum of $\mathrm{Clip}(D,\tau):=\big{(}\mathrm{Clip}(x_{i},\tau)\mid i=1,\dots,n\big{)}$ and add a Laplace noise of scale $O(\tau/\varepsilon)$. Note that the clipping introduces a (negative) bias of magnitude up to $\mathrm{Max}(D)\cdot|\{i\in[n]\mid x_{i}>\tau\}|$, where $\mathrm{Max}(D):=\max_{i}x_{i}$. Thus, one should choose a good clipping threshold $\tau$ that balances the DP noise and bias. Importantly, this must be done in a DP fashion, and this is where all past investigations on the clipping mechanism have been devoted. In the central model, the best error bound achievable is $O\big{(}\mathrm{Max}(D)\cdot\log\log U/\varepsilon\big{)}$ [DY23]. For the real summation problem, such an error bound is considered (nearly) instance-optimal, since any DP mechanism has to incur an error of $\Omega(\mathrm{Max}(D))$ on either $D$ or $D-\{\mathrm{Max}(D)\}$ [Vad17]. It has also been shown that the optimality ratio $O(\log\log U/\varepsilon)$ is the best possible [DY23].
@@
As suggested in [HLY21], the clipping mechanism can be easily implemented in the shuffle model as well, but requiring two rounds. The first round finds $\tau$. Then we broadcast $\tau$ to all users. In the second round, we invoke a summation protocol, e.g., the one in [${}^{+}$21], on $\mathrm{Clip}(D,\tau)$. Two-round protocols are generally undesirable, not only because of the extra latency and coordination overhead, but also because they leak some information to the users ($\tau$ in this case, which is an approximation of $\mathrm{Max}(D)$). Note that the shuffle model, in its strict definition, only allows one-way messages from users to the analyzer (through the shuffler), so the users should learn nothing from each other. Moreover, the central-DP mechanism for finding the optimal $\tau$ [DY23] does not work in the shuffle model. Instead, [HLY21] uses the complicated range-counting protocol of [${}^{+}$21]. This results in a sum estimation protocol that runs in two rounds, having an error of $\tilde{O}(\mathrm{Max}(D)\cdot\log^{3.5}U\sqrt{\log(1/\delta)})$ [^2] For any function $f$, $\tilde{O}(f):=f\cdot\mathrm{polylog}(f)$., and sends $\tilde{O}\big{(}\log^{6}U\log(1/\delta)/\varepsilon\big{)}$ messages per user. Thus, this protocol is of only theoretical interest; indeed, no experimental results are provided in [HLY21].
@@ 1.1 Our results
$1$-D Sum Prior work [${}^{+}$21] $O\big{(}U/\varepsilon\big{)}$ $1+o(1)$
@@
[HLY21] + [${}^{+}$21] + [${}^{+}$21] $\tilde{O}\big{(}\mathrm{Max}(D)\cdot\log^{3.5}U$ $\sqrt{\log(1/\delta)}/\varepsilon\big{)}$ Round $1$: $\tilde{O}\big{(}\log^{6}U\cdot\log\big{(}1/\delta\big{)}/\varepsilon\big{)}$ Round $2$: $1+o(1)$
@@
[HLY21] + [${}^{+}$21] ${}^{+}$21] $\tilde{O}\big{(}\mathrm{Max}_{\ell_{2}}(D)\cdot\big{(}\sqrt{d\log(nd)\log(1/\delta)}$ $+\log^{3.5}U_{\ell_{2}}\cdot\sqrt{\log(1/\delta)}\big{)}/\varepsilon\big{)}$ Round $1$: $\tilde{O}\big{(}\log^{6}U_{\ell_{2}}\cdot\log(1/\delta)/\varepsilon\big{)}$ Round $2$: $d+\tilde{O}(d^{1.5}\log^{1.5}(1/\delta)/(\varepsilon\sqrt{n})\big{)}$
@@
Table 1: Comparison between our results and prior works on sum estimation, high-dimensional sum estimation, and sparse vector aggregation under shuffle model, where we use the absolute error, $\ell_{2}$ error, and $\ell_{\infty}$ error as the error metrics. Each message contains $O(\log U+\log d+\log n)$ bits. The mechanism without an indicator for the round runs in a single round.
@@ 1.1.1 Contributions
(1) Sum estimation. For the vanilla sum estimation problem [^3] Note that although we focus on the integer domain $\{0,\dots,U\}$, our protocol easily extends to the real summation problem, where each value $x_{i}$ is a real number from $[0,1]$, by discretizing $[0,1]$ into $U$ buckets of width $1/U$. This incurs an extra additive error of $O(n/U)$. Thanks to the double logarithmic dependency on $U$, we could set $U$ sufficiently large (e.g., $U=n^{\log n}$) to make this additive error negligible while keeping the $O(\mathrm{Max}(D)\cdot\log\log n/\varepsilon)$ error bound. , we present a single-round protocol (Section 4 and 5) that achieves the optimal error of $O(\mathrm{Max}(D)\cdot\log\log U/\varepsilon)$ while sending $1+o(1)$ messages per client. This improves the error rate of $O(U/\varepsilon)$ from [${}^{+}$21] exponentially in $U$.
@@ 2 Related Work
For sum estimation under central-DP, the worst-case optimal error $O(U/\varepsilon)$ can be easily achieved by the Laplace mechanism. Many papers have studied how to obtain instance-specific error, i.e., an error depending on $\mathrm{Max}(D)$ [AD20, ATMR21, ${}^{+}$19, MRTZ17, HLY21, ${}^{+}$22, FDY22, DKSS23, DY23]. Most of these works rely on the clipping mechanism [AD20, ATMR21, ${}^{+}$19, MRTZ17, HLY21, ${}^{+}$22, DKSS23, DY23]. Similarly, for the high-dimensional sum aggregation problem, existing approaches have achieved instance-specific error by using the clipping mechanism [KLSU19, KSU20, BDKU20, HLY21, DSY23]; such mechanisms also yield an $\ell_{\infty}$ error of $\mathrm{Max}_{\ell_{2}}(D)$ for sparse vector aggregation.
In the shuffle-DP setting, for sum estimation, two settings are used. In the single-message setting, each user sends one message. Here, [BBGN19] achieve an error of $O(Un^{1/6})$ and further show that this is worst-case optimal. In the multi-message setting, where each user is allowed to send multiple messages, most prior works try to achieve the worst-case optimal error while minimizing communication costs. Cheu et al. [${}^{+}$19] first achieved an error of $O(U\sqrt{\log(1/\delta)}/\varepsilon)$ with $O(\sqrt{n})$ messages sent per user. Then, [GPV19] achieved the same error but reduced the number of messages per user to $O(\log(n))$. [GMPV20] and [BBGN20] further improved the error to $O(U/\varepsilon)$ with constant messages. Recently, [${}^{+}$21] reduced that communication to $1+o(1)$ messages per user. We aim to obtain instance-optimal error, while keeping the $1+o(1)$ per-client message complexity of [${}^{+}$21].
@@ 15       Add $z_{j}$ copies of $j$ into $S_{i}$;
Algorithm 1 Randomizer of $\mathrm{BaseSumDP}$ [${}^{+}$21].
@@ 4       $\widetilde{\mathrm{Sum}}(D)\leftarrow\widetilde{\mathrm{Sum}}(D)\cdot B$;
Algorithm 2 Analyzer of $\mathrm{BaseSumDP}$ [${}^{+}$21].

### v2

# Almost Instance-optimal Clipping for Summation Problems in the Shuffle Model of Differential Privacy

@@ Abstract
Differentially private mechanisms achieving worst-case optimal error bounds (e.g., the classical Laplace mechanism) are well-studied in the literature. However, when typical data are far from the worst case, instance-specific error bounds—which depend on the largest value in the dataset—are more meaningful. For example, consider the sum estimation problem, where each user has an integer $x_{i}$ from the domain $\{0,1,\dots,U\}$ and we wish to estimate $\sum_{i}x_{i}$. This has a worst-case optimal error of $O(U/\varepsilon)$, while recent work has shown that the clipping mechanism can achieve an instance-optimal error of $O(\max_{i}x_{i}\cdot\log\log U/\varepsilon)$. Under the shuffle model, known instance-optimal protocols are less communication-efficient. The clipping mechanism also works in the shuffle model, but requires two rounds: Round one finds the clipping threshold, and round two does the clipping and computes the noisy sum of the clipped data. In this paper, we show how these two seemingly sequential steps can be done simultaneously in one round using just $1+o(1)$ messages per user, while maintaining the instance-optimal error bound. We also extend our technique to the high-dimensional sum estimation problem and sparse vector aggregation (a.k.a. frequency estimation under user-level differential privacy).
@@ 1 Introduction
The shuffle model [BEM+17, CSU+19, EFM+19] of differential privacy (DP) is widely-studied in the context of DP computation over distributed data. The model has 3 steps: (1) Each client uses a randomizer $\mathcal{R}(\cdot)$ to privatize their data. (2) A trusted shuffler randomly permutes the inputs from each client and passes them to an untrusted analyzer, which (3) conducts further analysis. Unlike the central model of DP, where a trusted curator has access to all the data, the shuffle model provides stronger privacy protection by removing the dependency on a trusted curator. Unlike the local model, where clients send noisy results to the analyzer directly, the addition of the trusted shuffler allows for a significantly improved privacy-accuracy tradeoff. For problems like bit counting, shuffle-DP achieves an error of $O(1/\varepsilon)$ with constant probability [^1] In Section 1, all stated error guarantees hold with constant probability. We will make the dependency on the failure probability $\beta$ more explicit in later sections. [GKMP20], matching the best error of central-DP, while the smallest error achievable under local-DP is $O(\sqrt{n}/\varepsilon)$ [BIK+17, DNOR16].
@@
The summation problem, a fundamental problem with applications in statistics [KLSU19, BDKU20, HLY21], data analytics [DFY+22, THMR20], and machine learning such as the training of deep learning models [SCS13, BST14, ACG+16, ASY+18] and clustering algorithms [Ste21, SK18], has been studied in many works under the shuffle model [BBGN19, CSU+19, GPV19, GMPV20, BBGN20, GKM+21]. In this problem, each user $i\in[n]:=\{1,\dots,n\}$ holds an integer $x_{i}\in\{0,1,\dots,U\}$ and the goal is to estimate $\mathrm{Sum}(D):=\sum_{i}x_{i}$, where $D=(x_{1},\dots,x_{n})$. All of these works for sum estimation under shuffle-DP focus on achieving an error of $O(U/\varepsilon)$. Such an error can be easily achieved under central-DP, where the curator releases the true sum after masking it with a Laplace noise of scale $U/\varepsilon$. In the shuffle-DP model, besides error, another criterion that should be considered is the communication cost. The recent shuffle-DP summation protocol of [GKM+21] both matches the error of central-DP and achieves optimal communication. More precisely, it achieves an error that is just $1+o(1)$ times that of the Laplace mechanism, while each user just sends in expectation $1+o(1)$ messages, each of a logarithmic number of bits.
@@
However, in real applications (as well as in [GKM+21]), $U$ must be set independently of the dataset; to account for all possible datasets, it should be conservatively large. For instance, if we only know that the $x_{i}$’s are 32-bit integers, then $U=2^{32}-1$. Then the error of $O(U/\varepsilon)$ could dwarf the true sum for most datasets. Notice that sometimes some prior knowledge is available, and then a smaller $U$ could be used. For example, if we know that the $x_{i}$’s are people’s incomes, then we may set $U$ as that of the richest person in the world. Such a $U$ is still too large for most datasets as such a rich person seldom appears in most datasets.
@@
Instance-Awareness.  The earlier error bound of $O(U/\varepsilon)$ can be shown to be optimal, but only in the worst case. When typical input data are much smaller than $U$, an instance-specific mechanism (and error bound) can be obtained—i.e., a mechanism whose error depends on the largest element of the dataset. In the example of incomes above, an instance-aware mechanism would achieve an error proportional to the actual maximum income in the dataset. This insight has recently been explored under the central model of DP [ATMR21, PSY+19, MRTZ17, HLY21, FDY22, DFY+22].
@@
A widely used technique for achieving instance-specific error bounds under central-DP is the clipping mechanism [ATMR21, PSY+19, MRTZ17, HLY21]. For some $\tau$, each $x_{i}$ is clipped to $\mathrm{Clip}(x_{i},\tau):=\min(x_{i},\tau)$. Then we compute the sum of $\mathrm{Clip}(D,\tau):=\big{(}\mathrm{Clip}(x_{i},\tau)\mid i=1,\dots,n\big{)}$ and add a Laplace noise of scale $O(\tau/\varepsilon)$. Note that the clipping introduces a (negative) bias of magnitude up to $\mathrm{Max}(D)\cdot|\{i\in[n]\mid x_{i}>\tau\}|$, where $\mathrm{Max}(D):=\max_{i}x_{i}$. Thus, one should choose a good clipping threshold $\tau$ that balances the DP noise and bias. Importantly, this must be done in a DP fashion, and this is where all past investigations on the clipping mechanism have been devoted. In the central model, the best error bound achievable is $O\big{(}\mathrm{Max}(D)\cdot\log\log U/\varepsilon\big{)}$ [DY23].222[DY23] achieves an error of $O(\mathrm{Max}(D)\log\log(\mathrm{Max}(D)))$ rather than the cited $O(\mathrm{Max}(D)\log\log U))$. The $\log\log(\mathrm{Max}(D))$ result is more meaningful for the unbounded domain setting where $U=\infty$, but in the shuffle-DP model, there is currently no known method can handle the unbounded domain case for any problem, including sum estimation. Our proposed mechanism also only supports the bounded domain case. Therefore, for simplicity, we ignored this minor difference and just cited the $\log\log U$ result. For the real summation problem, such an error bound is considered (nearly) instance-optimal, since any DP mechanism has to incur an error of $\Omega(\mathrm{Max}(D))$ on either $D$ or $D-\{\mathrm{Max}(D)\}$ [Vad17]. The factor of $\log\log U/\varepsilon$ is known as the “optimality ratio”. It has been shown that the optimality ratio $O(\log\log U/\varepsilon)$ is the best possible in the case of $\delta=0$ ($\delta$ is a privacy parameter, see Section 3.1 for more details) [DY23]. Under the case of $\delta>0$, it is still an open question whether that ratio is optimal or not. Notably, in the literature [AD20, DFY+22, HLY21], a polylog optimality ratio is often considered satisfactory enough and is named as “instance-optimal” and so far no known DP mechanism has a better optimality ratio.
@@
As suggested in [HLY21], the clipping mechanism can be easily implemented in the shuffle model as well, but requiring two rounds. The first round finds $\tau$. Then we broadcast $\tau$ to all users. In the second round, we invoke a summation protocol, e.g., the one in [GKM+21], on $\mathrm{Clip}(D,\tau)$. Two-round protocols are generally undesirable, not only because of the extra latency and coordination overhead, but also because they leak some information to the users ($\tau$ in this case, which is an approximation of $\mathrm{Max}(D)$). Note that the shuffle model, in its strict definition, only allows one-way messages from users to the analyzer (through the shuffler), so the users should learn nothing from each other. Moreover, the central-DP mechanism for finding the optimal $\tau$ [DY23] does not work in the shuffle model. Instead, [HLY21] uses the complicated range-counting protocol of [GGK+21]. This results in a sum estimation protocol that runs in two rounds, having an error of $\tilde{O}(\mathrm{Max}(D)\cdot\log^{3.5}U\sqrt{\log(1/\delta)})$ [^3] For any function $f$, $\tilde{O}(f):=f\cdot\mathrm{polylog}(f)$., and sends $\tilde{O}\big{(}\log^{6}U\log(1/\delta)/\varepsilon\big{)}$ messages per user. Thus, this protocol is of only theoretical interest; indeed, no experimental results are provided in [HLY21].
@@ 1.1 Our results
$1$-D Sum Prior work [GKM+21] $O\big{(}U/\varepsilon\big{)}$ $1+o(1)$
@@
[HLY21] + [GKM+21] + [GGK+21] $\tilde{O}\big{(}\mathrm{Max}(D)\cdot\log^{3.5}U$ $\sqrt{\log(1/\delta)}/\varepsilon\big{)}$ Round $1$: $\tilde{O}\big{(}\log^{6}U\cdot\log\big{(}1/\delta\big{)}/\varepsilon\big{)}$ Round $2$: $1+o(1)$
@@
[HLY21] + [GKM+21] +[GGK+21] $\tilde{O}\big{(}\mathrm{Max}_{\ell_{2}}(D)\cdot\big{(}\sqrt{d\log(nd)\log(1/\delta)}$ $+\log^{3.5}U_{\ell_{2}}\cdot\sqrt{\log(1/\delta)}\big{)}/\varepsilon\big{)}$ Round $1$: $\tilde{O}\big{(}\log^{6}U_{\ell_{2}}\cdot\log(1/\delta)/\varepsilon\big{)}$ Round $2$: $d+\tilde{O}(d^{1.5}\log^{1.5}(1/\delta)/(\varepsilon\sqrt{n})\big{)}$
@@
Table 1: Comparison between our results and prior works on sum estimation, high-dimensional sum estimation, and sparse vector aggregation under shuffle model, where we use the absolute error, $\ell_{2}$ error, and $\ell_{\infty}$ error as the error metrics. Each message contains $O(\log U+\log d+\log n)$ bits. The mechanism without an indicator for the round runs in a single round. Our communication cost for 1-D Sum requires the condition $n=\omega(\log^{2}U)$.
@@ 1.1.1 Contributions
(1) Sum estimation. For the vanilla sum estimation problem [^4] Note that although we focus on the integer domain $\{0,\dots,U\}$, our protocol easily extends to the real summation problem, where each value $x_{i}$ is a real number from $[0,1]$, by discretizing $[0,1]$ into $U$ buckets of width $1/U$. This incurs an extra additive error of $O(n/U)$. Thanks to the double logarithmic dependency on $U$, we could set $U$ sufficiently large (e.g., $U=n^{\log n}$) to make this additive error negligible while keeping the $O(\mathrm{Max}(D)\cdot\log\log n/\varepsilon)$ error bound. , we present a single-round protocol (Section 4 and 5) that achieves the optimal error of $O(\mathrm{Max}(D)\cdot\log\log U/\varepsilon)$, which improves the error rate of $O(U/\varepsilon)$ from [GKM+21] exponentially in $U$. More importantly, we have $1+o(1)$ messages per client when $n=\omega(\log^{2}U)$ (see Theorem 5.1 for more details), a criterion typically met in most common regimes.
@@ 2 Related Work
For sum estimation under central-DP, the worst-case optimal error $O(U/\varepsilon)$ can be easily achieved by the Laplace mechanism. Many papers have studied how to obtain instance-specific error, i.e., an error depending on $\mathrm{Max}(D)$ [AD20, ATMR21, PSY+19, MRTZ17, HLY21, DFY+22, FDY22, DKSS23, DY23]. Most of these works rely on the clipping mechanism [AD20, ATMR21, PSY+19, MRTZ17, HLY21, DFY+22, DKSS23, DY23]. Similarly, for the high-dimensional sum aggregation problem, existing approaches have achieved instance-specific error by using the clipping mechanism [KLSU19, KSU20, BDKU20, HLY21, DSY23]; such mechanisms also yield an $\ell_{\infty}$ error of $\mathrm{Max}_{\ell_{2}}(D)$ for sparse vector aggregation.
In the shuffle-DP setting, for sum estimation, two settings are used. In the single-message setting, each user sends one message. Here, [BBGN19] achieve an error of $O(Un^{1/6})$ and further show that this is worst-case optimal. In the multi-message setting, where each user is allowed to send multiple messages, most prior works try to achieve the worst-case optimal error while minimizing communication costs. Cheu et al. [CSU+19] first achieved an error of $O(U\sqrt{\log(1/\delta)}/\varepsilon)$ with $O(\sqrt{n})$ messages sent per user. Then, [GPV19] achieved the same error but reduced the number of messages per user to $O(\log(n))$. [GMPV20] and [BBGN20] further improved the error to $O(U/\varepsilon)$ with constant messages. Recently, [GKM+21] reduced that communication to $1+o(1)$ messages per user. We aim to obtain instance-optimal error, while keeping the $1+o(1)$ per-client message complexity of [GKM+21].
@@ 15       Add $z_{j}$ copies of $j$ into $S_{i}$;
Algorithm 1 Randomizer of $\mathrm{BaseSumDP}$ [GKM+21].
@@ 4       $\widetilde{\mathrm{Sum}}(D)\leftarrow\widetilde{\mathrm{Sum}}(D)\cdot B$;
Algorithm 2 Analyzer of $\mathrm{BaseSumDP}$ [GKM+21].